# Exceptions to the Erdős–Straus–Schinzel conjecture

### v1

@@ Abstract.
A famous conjecture of Erdős and Straus is that for every integer $n\geq 2$, $4/n$ can be represented as $1/x+1/y+1/z$, where $x,y,z$ are positive integers. This conjecture was generalized to $5/n$ by Sierpiński, and then Schinzel conjectured that for every integer $m\geq 4$ there is a bound $n_{m}$ such that the fraction $m/n$ is the sum of 3 unit fractions for all integers $n\geq n_{m}$. Leveraging and generalizing work of Elsholtz and Tao, we show that if $n_{m}$ exists it must be at least $\exp(m^{1/3+o(1)})$; that is, there are numbers $n$ this large for which $m/n$ is not the sum of 3 unit fractions. We prove a weaker, but numerically explicit version of this theorem, showing that for $m\geq 6.52\times 10^{9}$ there is a prime $p\in(m^{2},2m^{2})$ with $m/p$ not the sum of 3 unit fractions, and report on some extensive numerical calculations that support this assertion with the much smaller bound $m\geq 19$. In addition we generalize a result of Vaughan to show that for each $m$, most $n$’s have $m/n$ representable, and we prove a result generalizing the problem to the sum of $j$ unit fractions.
@@ 1. Introduction
Egyptian fractions have a long and colorful history. According to the Rhind Papyrus (ca. 1550 BCE), ancient Egyptians preferred to write fractions as sums of unit fractions (fractions with numerator 1). We have not seen a compelling argument for why they had this preference, but nevertheless it opened the door to many intriguing problems. For surveys of some of the many problems and results, see [4, Ch. 4], [5, Sec. D11].
The Erdős–Straus conjecture, which dates to around 1948, asserts that $4/n$ is the sum of 3 unit fractions for every integer $n\geq 2$. An early result of Obláth [10] is that $n$ has this property if $n+1$ is divisible by a prime $p\equiv 3\pmod{4}$. This implies that asymptotically all $n$ have the Erdős–Straus property, the number of possible exceptions up to $N$ being $O(N/\sqrt{\log N})$. The count of possible exceptions has been strongly improved, though not recently: In 1970, Vaughan [12] gave the upper bound $N/\exp(c(\log N)^{2/3})$ for a positive constant $c$.
@@
There is an absolute positive constant $C$ such that for each pair $m,N$ with $4\leq m\leq(\log N)^{2}$ the number of $n\leq N$ with $m/n$ not the sum of $3$ unit fractions is at most $N/\exp(C(\log^{2}(N)/\varphi(m))^{1/3})$.
@@ 4. An upper bound
where $\tau_{3}(n)$ is the number of triples $a,b,c$ of integers with $abc=n$ (cf. Koukoulopoulos [6, Ex. 20.2]). Then for a prime $p\equiv-1\pmod{m}$,

### v2

@@ Abstract.
A famous conjecture of Erdős and Straus is that for every integer $n\geq 2$, $4/n$ can be represented as $1/x+1/y+1/z$, where $x,y,z$ are positive integers. This conjecture was generalized to $5/n$ by Sierpiński, and then Schinzel conjectured that for every integer $m\geq 4$ there is a bound $n_{m}$ such that the fraction $m/n$ is the sum of 3 unit fractions for all integers $n\geq n_{m}$. Leveraging and generalizing work of Elsholtz and Tao, we show that if $n_{m}$ exists it must be at least $\exp(m^{1/3+o(1)})$; that is, there are numbers $n$ this large for which $m/n$ is not the sum of 3 unit fractions. We prove a weaker, but numerically explicit version of this theorem, showing that for $m\geq 6.52\times 10^{9}$ there is a prime $p\in(m^{2},2m^{2})$ with $m/p$ not the sum of 3 unit fractions, and report on some extensive numerical calculations that support this assertion with the much smaller bound $m\geq 20$. A result of Vaughan is that for each $m$, most $n$’s have $m/n$ representable; we make the dependence on $m$ in this result explicit. In addition, we prove a result generalizing the problem to the sum of $j$ unit fractions.
@@ 1. Introduction
Egyptian fractions have a long and colorful history. According to the Rhind Papyrus (ca. 1550 BCE), ancient Egyptians preferred to write fractions as sums of unit fractions (fractions with numerator 1). We have not seen a compelling argument for why they had this preference, but nevertheless it opened the door to many intriguing problems. For surveys of some of the many problems and results, see [4, Ch. 4], [6, Sec. D11].
The Erdős–Straus conjecture, which dates to around 1948, asserts that $4/n$ is the sum of 3 unit fractions for every integer $n\geq 2$. An early result of Obláth [11] is that $n$ has this property if $n+1$ is divisible by a prime $p\equiv 3\pmod{4}$. This implies that asymptotically all $n$ have the Erdős–Straus property. In fact, exceptions should only be divisible by primes $p$ for which $p+1$ has no prime factor $q\equiv 3\pmod{4}$. The number of such primes $p\leq N$ is $O(N/\log^{3/2}N)$ as can be seen from sieve methods. This property allows only $O(N/\log^{3/2}N)$ exceptional numbers up to $N$ via an argument akin to the Hardy–Ramanujan inequality (cf. Gottschlich [5, Lemma 2.3]). The count of possible exceptions has been strongly improved, though not recently: In 1970, Vaughan [13] gave the upper bound $N/\exp(c\log^{2/3}\kern-2.0ptN)$ for a positive constant $c$.
@@
There is an absolute positive constant $C$ such that for each pair $m,N$ with $4\leq m\leq\log^{2}N$ the number of $n\leq N$ with $m/n$ not the sum of $3$ unit fractions is at most $N/\exp(C\log^{2/3}(N)/\varphi(m)^{1/3})$.
@@ 4. An upper bound
where $\tau_{3}(n)$ is the number of triples $a,b,c$ of integers with $abc=n$ (cf. Koukoulopoulos [7, Ex. 20.2]). Then for a prime $p\equiv-1\pmod{m}$,
@@ 10. Some final thoughts
Acknowledgment
Thanks are due to Paul Pollack for some helpful comments.